The sum of 3 numbers is 94, if the first number is 3/5 of the second number and the second number is 4/3 of the third. What are the three numbers?
Answer by stanbon(75887)

The sum of 3 numbers is 94, if the first number is 3/5 of the second number and the second number is 4/3 of the third. What are the three numbers?
------
Equations:
x + y + z = 94
x = (3/5)y
y = (4/3)z
-----
Substitute for "x" and "Z" and solve for "y":
----
(3/5)y + y + (3/4)y = 94
----
Multiply thru by 20 to get:
12y + 20y + 15y = 94*20
47y = 94*20
y = 2*20
y = 40
-----
x = (3/5)y = 24
z = (3/4)y = 30
=====================
Cheers,
Stan H.
=====================